order: dobutrex 3mcg/kg/min
available: Dobutrex 500mg in 250mL D5 1/2 NS
clients weight: 149 lb
the IV flow rate is
Answer #1

ANSWER.

* Calculation of IV flow rate.

       Flow rate(ml/hr) = Volume(ml) /Time(hr).

Calculation of volume(ml).

          Formula = Desired dose/Dose in hand * Quantity.

Here

    Ordered dose = 3 mcg/kg/min.

Convert mcg to mg for calculation

           1 mcg = 0.001 mg.

   Ordered dose = 3 * 0.001

                          = 0.003 mg/kg/min.

Weight of the patient = 149 lbs.

    Convert lbs to kg for calculation

               1 lbs = 0.454 kg.

    Weight of the patient = 149 * 0.454

                                       = 67.6 kg.

   Desired dose (D) = 0.003 * 67.6

                                      = 0.20 mg/min.

         Dose in hand (H) = 500 mg.

             Quantity(Q) = 250 ml.

As per above formula

            Volume(ml) of drug required = D/H * Q

                                                          = 0.20/500 * 250

                                                          = 0.0004 * 250

                                                          = 0.1 ml/min.

       Convert ml/min to ml/hr multiply with 60.

                     ml/hr = 0.1 * 60

                               = 6 ml/hr.